OpenStudy (goformit100):

A radar station locates a sinking ship at range 17.3 km and bearing 136° clockwise from north. From the same station a rescue plane is at horizontal range 19.6 km, 153° clockwise from north, with elevation 2.20 km. (a) Write the vector displacement from plane to ship, letting i represent east, j north, and k up. (b) How far apart are the plane and ship?

OpenStudy (goformit100):

@mikaela19900630

OpenStudy (anonymous):

for a) Converting to CCW from + x axis 360-(136-90) = 314 360-(153-90) = 297 Position vector of ship relative to origin: 17.3* Cos( 314) i+ 17.3* Sin( 314) j or Rs = <12.02 i , -12.44 j , 0k> Position vector of plane relative to origin "below" : 19.6* Cos( 297) i+ 19.6* Sin( 297) j or Rp = <8.90 i , -17.46 j , 2.2 k> Rsp = Rso - Rpo = 12.02 - 8.90 = 3.12i km "east" -12.44 - (-17.46) = 5.02j km "north" 0 - 2.2 = -2.2k km "below" The plane sees the ship at: Rsp =<3.12i , 5.02j , -2.2k> km

OpenStudy (anonymous):

For b) |R| = sqr(3.12^2+(5.02)^2 + (-2.2)^2) = 6.307 km